PULSED-FIELD IONIZATION ELECTRON AND PHOTOIONIZATION SPECTROSCOPY OF Li-NAPHTHALENE AND ANTHRACENE COMPLEXES

\underline{SERGIY KRASNOKUTSKI}, JUNG SUP LEE, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Li-naphthalene (C_{10}H_{8}) and anthracene (C_{14}H_{10}) were produced by reactions of laser-vaporized Li atoms and ligand vapor. Although naphthalene has a sufficient vapor pressure at the room temperature, the vapor pressure of anthracene is rather low. Therefore, anthracene was co-vaporized by laser ablation of a target made of Li and organic powders. The pulsed-field ionization electron spectrum of Li-naphthalene shows a band origin at 35489 (5) cm^{−1} and extensive vibrational structures with intervals in the 140-1400 cm^{−1} range. These vibrational structures were assigned to excitations of Li-naphthalene stretches and bends as well as in-plane and out-of-plane ring deformations. This complex has a Cs structure with Li located above one of the rings, and the naphthalene plane is significantly distorted in the neutral state. Unlike naphthalene where two six-membered rings are the same, the side and central rings are different in anthracene. Although vibrationally resolved electron spectrum has not been obtained, the photoionization efficiency spectrum of Lianthracene exhibits two thresholds at 33800 (300) and 37200 (200) cm^{−1}, which correspond to ionization of two structural isomers with Li binding to the side and central rings, respectively.